Circular No. 8317

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004be AND 2004bf Further to IAUC 8312, K. Shimasaki, M. Moore, and W. Li report the

LOSS discovery of two apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004be Apr. 6.28 10^{h}00^{m}19.^{s}47 −24^{o}48′13′′.8 17.0 4′′.0 W, 2′′.4 N
2004bf Apr. 7.44 13 49 15.45 +35 15 12.5 17.5 17′′.1 E, 14′′.8 S

Additional KAIT magnitudes: SN 2004be in ESO 499-G34, Mar. 28.25 UT, [18.0; Apr. 8.22, 16.7. SN 2004bf in UGC 8739, Mar. 6.45, [18.5; 23.48, [18.0; Apr. 8.43, 17.5.

SUPERNOVA 2004bg IN UGC 6363 Further to IAUC 8316, M. Armstrong reports his discovery of an appar-

ent supernova (mag ∼ 15.5) on unfiltered CCD images taken on Apr. 7.965 and 8.936 UT. SN 2004bg is located at α = 11^{h}21^{m}01^{s}.53, δ = +21^{o}20′23′′.4 (equinox 2000.0), which is ≈ 14′′.2 east and 7′′.4 north of the center of UGC 6363. Nothing was visible at this position on Armstrong’s image taken on 2003 Mar. 7 or on Palomar Sky survey red and blue plates. M. Mobberley, Cockfield, U.K., confirmed the new object on an unfiltered CCD image taken on Apr. 8.828 with a 0.35-m reflector.

SUPERNOVA 2004bd IN NGC 3786 T. Hamane, K. Kinugasa, and H. Kawakita, Gunma Astronomical Ob-

servatory (GAO); and H. Yamaoka, Kyushu University, report that a lowresolution spectrum (range 460–660 nm) of SN 2004bd (cf. IAUC 8316), obtained with the GAO 0.65-m telescope (+ GCS) on Apr. 8.5 UT, shows it to be a type-Ia supernova near maximum. The expansion velocity, determined from the Si II 635-nm absorption minimum, is ∼ 9600 km/s.

COMET C/2001 Q4 (NEAT) Visual total-magnitude estimates: Mar. 17.40 UT, 6.3 (D. A. J. Sear-

gent, Cowra, N.S.W., 2×25 binoculars); 21.37, 6.3 (S. T. Rae, Tamahere, New Zealand, 8×21 binoculars); 25.00, 6.0 (J. Shanklin, Stanley, Falkland Islands, naked eye); 28.01, 5.6 (S. F. Hoenig, La Silla, Chile, naked eye); Apr. 1.76, 5.6 (Seargent, The Entrance, N.S.W., naked eye).

2004 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green